OpenStudy (anonymous):

Find the assessed value if the market value is $80,000 and the percent of assessment is 70%.

OpenStudy (anonymous):

(70/100)*80000 = 70 * 800 = 56000

OpenStudy (anonymous):

GBU too!

OpenStudy (anonymous):

Thank you! :D

OpenStudy (anonymous):

ur welcome?!